Write an equation in standard form of the line whose slope is 1/2 (one half) and whose y-intercept is -3.
Answer by jejess(7)

y=mx+b (standard form)
m= slope
b= y-intercept
y=1/2x+(-3)
y=1/2x-3
make sure you write down what you know from the question first so you don't get lost in the words.
You may have gotten confused because you didn't know what standrard form was, just remember when dealing with linear equations it is always y=mx+b
m is always slope and b is always the y-intercept.
This formula is easy to change around to slove for unknowns.